OpenStudy (anonymous):

[y^2-2y/y^2+7y-18] X [y^2-81/y^2-11y+18] Simplify the rational expression.

OpenStudy (anonymous):

break the denominators off into factored squares, and cancel terms to simplify.

OpenStudy (anonymous):

i did that and got y(y-9)/(x-2) (y+2)

OpenStudy (anonymous):

polpak will help you out

OpenStudy (anonymous):

\[\frac{y(y - 2)}{(y+9)(y-2)} * \frac{(y+9)(y-9)}{(y+9)(y+2)} \]

OpenStudy (anonymous):

oh wait a sec.....on the bottom right i put (y-2) (y+2) for some reason. the rest i got right though.

OpenStudy (anonymous):

hold on let me try this to see if i get it right

OpenStudy (anonymous):

Leaves \[\frac{y(y+9)}{(y+9)(y+2)} = \frac{y(y-9)}{(y+9)(y+2)}\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

polpak is riqht but i think u still have to cancel the terms

OpenStudy (anonymous):

he did

OpenStudy (anonymous):

ohhh yeah he did... i didnt see the equals siqn... sorry! :))

OpenStudy (anonymous):

[y^2-2y/y^2+7y-18] factor you get [y(y-2)/(y-2)(y+9)]cancel you get y/(y+9) [y^2-81/y^2-11y+18] factor you get [(y+9)(y-9)]/[(y-2)(y-9)] cancel you get (y+9)/(y-2) now you can time; [y/(y+9)]*[(y+9)/(y-2)] [y^2+9y)]*[(y^2+7y-14)] from here can you do your own?

OpenStudy (anonymous):

would you get rid of y-2??? cuz thats the only way im getting the y(y+9)

OpenStudy (anonymous):

you foil

OpenStudy (anonymous):

time

OpenStudy (anonymous):

[y/(y+9)]*[(y+9)/(y-2)] sorry time not divi [y/(y+9)]*[(y+9)*(y-2)]

OpenStudy (anonymous):

Whoops, was that a -11 in the denominator of the second fraction?

OpenStudy (anonymous):

My font makes it hard to read those sometimes. Lets try this again.

OpenStudy (anonymous):

okay :D ha.

OpenStudy (anonymous):

srry I wrong again,it hard to tye ,easy for me to write

OpenStudy (anonymous):

\[\frac{y^2-2y}{y^2+7y-18} \times \frac{y^2-81}{y^2-11y+18}\] \[ = \frac{y(y-2)}{(y+9)(y-2)} \times \frac{(y+9)(y-9)}{(y-9)(y-2)}\] \[ = \frac{y}{(y+9)} \times \frac{(y+9)}{(y-2)}\] \[ = \frac{y(y+9)}{(y+9)(y-2)} \] \[= \frac{y}{y-2}\]

OpenStudy (anonymous):

[y^2+9y)]*[(y^2+7y-14)]not this [y^2+9y)]*[(y^2+7y-18)] this correct

OpenStudy (anonymous):

While that's correct, yours can be simplified further Nancy.

OpenStudy (anonymous):

Well correct assuming that by * you mean / that is.

OpenStudy (anonymous):

ha okay thanks polpak and nancy i really appreciate your help!

OpenStudy (anonymous):

Did you follow all that Carr?

OpenStudy (anonymous):

The best strategy to use with these is to leave them factored as long as possible so you can quickly see what cancels.

OpenStudy (anonymous):

yeah i got a little head of my self and put the wrong number but thank you both!

OpenStudy (anonymous):

That's ok, I misread a negative so we all make mistakes ;)

OpenStudy (anonymous):

yes still time (y/y-2)*(y+9)(y-2)*[y^2-81/y^2-11y+18]

OpenStudy (anonymous):

from here are you OK to do your own? [y^2+9y)]*[(y^2+7y-18)]